# Modeling the Carbon Footprint of HPC: The Top 500 and EasyC

### v1

@@ Abstract
With the disclosed Top500.org data, and using a new tool, EasyC, we were able to model the operational carbon of 391 HPC systems and the embodied carbon of 283 HPC systems. We further show how this coverage can be enhanced by exploiting additional public information. With improved coverage, then interpolation is used to produce the first carbon footprint estimates of the Top 500 HPC systems. They are 1,393.7 million MT CO2e operational carbon (1 Year) and 1,881.8 million MT CO2e embodied carbon. We also project how the Top 500’s carbon footprint will increase through 2030.
@@ I Introduction
  • The first assessment of Top500 carbon footprint. Annual operational carbon is 1,393.7 million MT CO2e (equivalent to 325k gasoline-powered vehicles annual emissions). Embodied carbon is 1,881.8 million MT CO2e (equivalent to 439k gasoline-powered vehicles annual emissions).
@@ IV-B Constructing an Estimate of the Full Top 500
To fill the gaps of 100’s of systems, we interpolate the carbon footprint for the systems missing data using the average of the nearest 10 peers (5 lower and 5 higher) in the Top 500.333If the peers are also incomplete, we use the next closest peers. The data in Figure 7a represent the carbon footprint of HPC. On the left, we present the EasyC carbon footprint for all the systems covered by Top500.org plus other public information. On the right, we include those and add the interpolated systems. Continuing, we also compute the average operational and embodied carbon of systems in the systems with enough data (for EasyC) and for the full 500, completed by interpolation. This brings the Operational carbon for the Top 500 to 1,393.7 million MT CO2e, adding the missing 10 systems increased operational footprint by only 1.74%.
@@
Adding the missing 96 systems increased Embodied carbon to 1,881.8 million MT CO2e, an increase of 23.18%. This operational carbon is equal to one year’s emissions for 325,000 gasoline-powered vehicles or 3.5 billion vehicle miles. The embodied carbon is equivalent to one year’s emissions for 439,000 gasoline-powered vehicles or 4.8 billion passenger miles. Of course the embodied carbon is 1-time for the lifetime of the computer system so it would be smaller if annualized. In Figure 8b, the average operational and embodied carbon for Top 500 systems is presented. Each is millions of MT CO2e, comparable to that of thousands of homes.
In Figure 7, we report the aggregate carbon footprint for the Top 500, both with the systems for which we have sufficient data. For 490 systems, the total operational carbon is 1,369.9 Million MT CO2e, and for 404 systems the total embodied carbon is 1,527.7 Million MT CO2e.
@@
As a sensitivity study, we show the difference that adding public information made in Figure 9. For operational carbon, we can see that adding public information does not change the overall results significantly. For operational carbon, we see refinement to a narrower region average carbon intensity (ACI) of power, which can increase or decrease by as much as 77.5%. But even a number of these do create a large change, the total change for the entire Top 500 is only 2.85% (38 million MTCO2e). For embodied carbon, there are larger changes, mostly increasing the carbon footprint, but these have small contribution. The biggest change is due to large systems where no estimate was previously possible (not shown). Overall this produces and increase of 670.48 million MTCO2e, for an 78% change overall. The use of novel accelerators, not documented in public information, is the largest problem. Approximating these accelerators with mainstream GPUs produces systematic underestimates of silicon size.
@@ IV-C How Carbon Footprint will Change for the Top 500?
We project carbon footprint of the Top 500 based on these growth rates (see Figure 10) Operational carbon’s 10.3% annual growth reflects both increasing computing power represented in the Top 500 systems, but also the slowing benefits of Moore’s Law and the reality of post-Dennard scaling, as power consumption continues to increase rapidly. By 2030, Top 500’s operational carbon is nearly double that of 2024. Clearly architectural customization and accelerators is not enough to reduce power consumption.
@@
The projection shows a continued increase in perf/carbon, but at an annual rate of 0.2 PFlop/s million MT CO2e. While there is still progress, this is dramatically slower than what we experienced for decades under Dennard scaling, and parallelism scaling where we’d expect 2x performance / unit power every 18-month generation (plotted for comparison). Note that the current increase in performance / unit carbon is not sufficient to compensate for the rapid growth of use of computing. As the net growth of total carbon footprint of the Top 500 continues at 10.3%/annum (see Figure 10).

### v2

@@ Abstract
With the disclosed Top500.org data, and using a new tool, EasyC, we were able to model the operational carbon of 391 HPC systems and the embodied carbon of 283 HPC systems. We further show how this coverage can be enhanced by exploiting additional public information. With improved coverage, then interpolation is used to produce the first carbon footprint estimates of the Top 500 HPC systems. They are 1.4 million MT CO2e operational carbon (1 Year) and 1.9 million MT CO2e embodied carbon. We also project how the Top 500’s carbon footprint will increase through 2030.
@@ I Introduction
  • The first assessment of Top500 carbon footprint. Annual operational carbon is 1.39 million MT CO2e (equivalent to 325k gasoline-powered vehicles annual emissions). Embodied carbon is 1.88 million MT CO2e (equivalent to 439k gasoline-powered vehicles annual emissions).
@@ IV-B Constructing an Estimate of the Full Top 500
To fill the gaps of 100’s of systems, we interpolate the carbon footprint for the systems missing data using the average of the nearest 10 peers (5 lower and 5 higher) in the Top 500.333If the peers are also incomplete, we use the next closest peers. The data in Figure 7a represent the carbon footprint of HPC. On the left, we present the EasyC carbon footprint for all the systems covered by Top500.org plus other public information. On the right, we include those and add the interpolated systems. Continuing, we also compute the average operational and embodied carbon of systems in the systems with enough data (for EasyC) and for the full 500, completed by interpolation. This brings the Operational carbon for the Top 500 to 1.39 million MT CO2e, adding the missing 10 systems increased operational footprint by only 1.74%.
@@
Adding the missing 96 systems increased embodied carbon to 1.88 million MT CO2e, an increase of 23.18%. This operational carbon is equal to one year’s emissions for 325,000 gasoline-powered vehicles or 3.5 billion vehicle miles. The embodied carbon is equivalent to one year’s emissions for 439,000 gasoline-powered vehicles or 4.8 billion passenger miles. Of course the embodied carbon is 1-time for the lifetime of the computer system so it would be smaller if annualized. In Figure 8b, the average operational and embodied carbon for Top 500 systems is presented. Each is thousands of MT CO2e, comparable to that of thousands of homes.
In Figure 7, we report the aggregate carbon footprint for the Top 500, both with the systems for which we have sufficient data. For 490 systems, the total operational carbon is 1.37 Million MT CO2e, and for 404 systems the total embodied carbon is 1.53 Million MT CO2e.
@@
As a sensitivity study, we show the difference that adding public information made in Figure 9. For operational carbon, we can see that adding public information does not change the overall results significantly. For operational carbon, we see refinement to a narrower region average carbon intensity (ACI) of power, which can increase or decrease by as much as 77.5%. But even a number of these do create a large change, the total change for the entire Top 500 is only 2.85% (38 thousand MT CO2e). For embodied carbon, there are larger changes, mostly increasing the carbon footprint, but these have small contribution. The biggest change is due to large systems where no estimate was previously possible (not shown). Overall this produces and increase of 670.48 thousand MT CO2e, for an 78% change overall. The use of novel accelerators, not documented in public information, is the largest problem. Approximating these accelerators with mainstream GPUs produces systematic underestimates of silicon size.
@@ IV-C How Carbon Footprint will Change for the Top 500?
We project carbon footprint of the Top 500 based on these growth rates (see Figure 10). Operational carbon’s 10.3% annual growth reflects both increasing computing power represented in the Top 500 systems, but also the slowing benefits of Moore’s Law and the reality of post-Dennard scaling, as power consumption continues to increase rapidly. By 2030, Top 500’s operational carbon is nearly double that of 2024. Clearly architectural customization and accelerators is not enough to reduce power consumption.
@@
The projection shows a continued increase in perf/carbon, but at an annual rate of 0.2 PFlop/s per thousand MT CO2e. While there is still progress, this is dramatically slower than what we experienced for decades under Dennard scaling, and parallelism scaling where we’d expect 2x performance / unit power every 18-month generation (plotted for comparison). Note that the current increase in performance / unit carbon is not sufficient to compensate for the rapid growth in the use of computing. As the net growth of total carbon footprint of the Top 500 continues at 10.3%/annum (see Figure 10).